OpenStudy (loser66):

Statistics help, please. On lecture, I have \(MSE(\hat\theta)=E(\hat\theta -\theta)^2=Var(\hat\theta)+(B_{\theta}(\hat\theta))^2\) I don't get it. Please, explain me

OpenStudy (loser66):

@Zarkon \((\hat\theta -\theta)^2=\hat\theta^2-2\hat\theta\theta +\theta^2\) Hence expectation of it is \(=E(\hat\theta^2)-2E(\hat\theta\theta)+E(\theta ^2)\) \(E(\hat\theta^2)-(E(\hat\theta))^2+(E(\hat\theta))^2-(2E(\hat\theta\theta)+E(\theta ^2)\) the first two terms gives \(Var(\hat\theta)\) but the last three terms = \((B_\theta(\hat\theta))^2\) if \(\hat\theta \) and \(\theta\) independent. Am I right? but they are not independent, right?